# suppSupplementary References

### v1

@@ B.2 ShapeNet
Table 6: Surface reconstruction metrics on ShapeNet [47].
@@
Table 8: Distance function query for near surface region on ShapeNet [47].
@@
Figure 14: Sectional views of ShapeNet signed distance function reconstruction results. Car, Lamp A, and Lamp B are shown in Fig. 8, while Lamp C, Rifle, and Watercraft are presented in Fig. 9 and Fig. 18. By cross-verifying with them, our HotSpotachieves reconstructions with fewer extra boundaries, more regularized level sets, and accurate topologies.
@@
B.3 Complex Topology Reconstruction
Figure 15: Visualizations of Genus6, Kangaroo, and Voronoi Sphere from the Mehta et al [1]. dataset. We show the number of iterations used in training in the parenthesis. We achieve excellent results with only half the iterations, successfully capturing the correct topologies.
@@ B.3 Complex Topology Reconstruction
B.4 Surface Reconstruction Benchmark (SRB)
@@ Appendix C Relation to PHASE [44]
Theorem 2 in the PHASE paper states that to achieve minimal area, $w_{b}$ should converge to 00 as $\epsilon$ approaches an infinitesimal value. However, this interpretation does not align with the nature of the signed distance function reconstruction task which should complete the manifold and connect points, as discussed in Section 3, and may lead to unintended consequences as follows: When the boundary weight $w_{b}$ is overly small, the signed distance function values of boundary points cannot even converge to close to 0. When the weight $w_{b}$ is still not large enough, their output collapses from a surface-based-distance signed distance function to a point-cloud-based-distance unsigned distance function upon reaching the target where the boundaries are only point clouds and the area becomes almost zero. In contrast, our model demonstrates robust and accurate performance, effectively connecting points and interpolating boundaries by taking advantages of neural network’s spectral bias [51], with a large boundary weight $w_{b}$ and an adaptive absorption $\lambda$, as discussed in Section 6.
@@
For their proposed setting $\epsilon=0.01$, only to get $s=2$, we will need to set $o=1-e^{-20}\approx 0.9999999979$, which already is beyond what a 32-bit floating point number can reliably represent. Rescaling the scene is unfortunately not going to help, since the parameter $\epsilon$ is scene dependent and needs to be scaled accordingly. In practice, we verify that when PHASE struggles with queries that are far away from the surfaces, and show visual examples in the inset, where the first figure is PHASE result with $\epsilon=0.01$ and occupancy clamped at $0.99$, and the second figure is our result, which can represent arbitrary signed distance function values.

### v2

@@ B.3 ShapeNet
Table 7: Surface reconstruction metrics on ShapeNet [55]. Bold and underlined data: optimal; bold only: suboptimal. Same below.
@@
Table 9: Distance function query metrics for near surface region on ShapeNet [55].
@@
Figure 16: Sectional views of ShapeNet [55] signed distance function reconstruction results. Car, Lamp A, and Lamp B are shown in Fig. 9, while Lamp C, Rifle, and Watercraft are presented in Fig. 10 and Fig. 20. By cross-verifying with them, HotSpot achieves reconstructions with fewer extra boundaries, more regularized level sets, and accurate topologies.
@@
B.4 Complex Topology Reconstruction
Figure 17: Visualizations of Genus6, Kangaroo, and Voronoi Sphere from Mehta et al [1]. We show the number of iterations used in training in the parentheses. We achieve excellent results with only half the iterations, successfully capturing the correct topologies.
@@ B.4 Complex Topology Reconstruction
B.5 Surface Reconstruction Benchmark (SRB)
@@ Appendix C Relation to PHASE [49]
Theorem 2 in the PHASE paper states that to achieve minimal area, $w_{b}$ should converge to 00 as $\epsilon$ approaches an infinitesimal value. However, this interpretation does not align with the nature of the signed distance function reconstruction task which should complete the manifold and connect points, as discussed at the end of Section 3, and may lead to unintended consequences as follows: When the boundary weight $w_{b}$ is overly small, the signed distance function values of boundary points cannot even converge to close to 0. When the weight $w_{b}$ is still not large enough, their output collapses from a surface-based-distance signed distance function to a point-cloud-based-distance unsigned distance function upon reaching the target where the boundaries are only point clouds and the area becomes almost zero. In contrast, our model demonstrates robust and accurate performance, effectively connecting points and interpolating boundaries by taking advantages of neural network’s spectral bias [58], with a large boundary weight $w_{b}$ and an adaptive absorption $\lambda$, as discussed in Section 6.
@@
For their proposed setting $\epsilon=0.01$, only to get $s=2$, we will need to set $o=1-e^{-20}\approx 0.9999999979$, which already is beyond what a 32-bit floating point number can reliably represent. Rescaling the scene is unfortunately not going to help, since the parameter $\epsilon$ is scene dependent and needs to be scaled accordingly.
In practice, we verify that when PHASE struggles with queries that are far away from the surfaces, and show visual examples in the inset, where the first figure is PHASE result with $\epsilon=0.01$ and occupancy clamped at $0.99$, and the second figure is our result, which can represent arbitrary signed distance function values.